If sin x = 0.20, cos x = .98, sin y = .12, and cos y = .99, which of the following is cos(x + y)?

A. -.94
B. -.95
C. .96
D. .95
1
Expert's answer
2014-03-07T05:24:30-0500
As cos(x+y)=cos(x)cos(y)-sin(x)sin(y) we obtain that cos(x+y)=0.99*0.98-0.2*0.12=0.9462. So, the answer is D.